COMPLETE LIST OF CONFERENCE TALKS, LECTURES,

AND OTHER INVITED PRESENTATIONS

Krzysztof Galicki

1. January 26, 2006 – Geometric Structures on 5-Manifolds, Ohio State University colloquium, Columbus, USA.

2. December 16, 2005 – Geometric Structures on 5-Manifolds, Universität Köln geometry seminar, Köln, Germany.

3. November 22, 2005 – Geometric Structures on 5-Manifolds, Humboldt Universität geometry seminar, Berlin, Germany.

4. November 14, 2005 – Geometric Structures on 5-Manifolds (in Polish), Jagiellonian University geometry seminar, Kraków, Poland.

5. November 1, 2005 – Seifert Bundles and Contact Structures, invited lecture at the Erwin Schrödinger Institute, Vienna, Austria.

6. October 10, 2005 – Manifolds with Positive Ricci Curvature, Arthur Besse seminar, École Polytechnique, Palaiseau, France.

7. September 23, 2005 – Sasakian Structures in Mathematics and Physics, invited lecture at the Géométrie Pseudo-Riemannienne et Supersymétrie, September 22-24, 2005; Institut de Recherche Mathématique Avancée de Strasbourg, Strasbourg, France.

8. September 1, 2005 – Geometric Structures on 5-manifolds, MPIM Oberseminar, MPIM, Bonn, Germany.

9. August 5, 2005 – Geometric Structures on 5-Manifolds, invited lecture at the SFB 478 workshop Curvature and Global Shape, 31 July – 7 August, 2005, Münster, Germany.

10. July 25, 2005 – Geometric Structures on 5-Manifolds, geometry seminar at the Università di Cagliari, Cagliari, Italy.

11. July 12, 2005 – Geometric Structures on 5-Manifolds, geometry seminar at the Università di Roma, “La Sapienza”, Rome, Italy.

12. July 3, 2005 – Geometric Structures on 5-manifolds, colloquium talk at Università di Bari, Bari, Italy.

13. June 7, 14, 2005 – Sasakian Geometry I-III, invited series of three lectures at the Institut Élie Cartan, Université Henri Poincaré, Nancy, France.

14. June 13, 2005 – Sasakian Geometry of 5-Manifolds, Arthur Besse seminar, École Polytechnique, Palaiseau, France.

15. December 10, 2004 – Einstein Metriken auf Exotischen Sphären (in German), Universität Hamburg, Hamburg, Germany.

16. November 19, 2004 – Einstein Metrics on Spheres, lecture given at the geometry seminar at the University of Bochum, Bochum, Germany.

17. November 2, 2004 – Einstein Metrics on Brieskorn Manifolds, lecture given at the geometry seminar at the University of Hamburg, Hamburg, Germany.

18. October 21, 2004 – Einstein Metrics on Spheres, lecture given at the geometry seminar at the University of Bonn, Bonn, Germany.

19. October 17, 2004 – Seifert Bundles, log del Pezzo Surfaces, and Einstein 5-Manifolds, MPI Oberseminar Topologie, Bonn, Germany.

20. October 12,14,15, 2004 – Sasakian Structures and Einstein Metrics I-III, series of three lectures at Centro di Ricerca Matematica Ennio De Giorgi, Differential Geometry and Topology, September 6 - November 12, 2004, Pisa, Italy.

21. May 12, 2004 – Transverse Fano Structures and Einstein Metrics on Spheres, invited colloquium, University of Bucharest, Bucharest, Romania.

22. April 22, 2004 – Transverse Fano Structures and Einstein Metrics on Spheres, invited colloquium, University of Göttingen, Göttingen, Germany.

23. April 6, 2004 – Transverse Fano Structures and Einstein Metrics on Spheres, invited lecture at the Erwin Schrödinger Institute in the in the program Geometric and analytic problems related to Cartan connections, Vienna, Austria.

24. March 18, 2004 – Transverse Fano Structures and Einstein Metrics on Spheres, MPIM Oberseminar, MPIM, Bonn, Germany.

25. February 28, 2004 – Einstein Metrics on Spheres, invited lecture at the Positive scalar curvature and related topics conference at LMU München, München, Germany.

26. February 2, 2004 – Transverse Fano Structures and Einstein Metrics on Spheres, geometry seminar at the University of Oxford, Oxford, England.

27. January 18-24, 2004 – Contact Structures, Complex Singularities, and Einstein Metrics I-III, a series of three invited plenary lectures at the XXIV Winter School in Geometry and Physics, Srni, Chech Republic.

28. July 21, 2003 – Transverse Fano Structures, Ricci Curvature, and Einstein Metrics, geometry seminar at the Università di Cagliari, Cagliari, Italy.

29. July 4, 2003 – Negative Self-Dual Einstein Metrics as Quotients, lecture given at Università di Roma, “La Sapienza”, Roma, Italy.

30. June 24, 2003 – Transverse Fano Structures, Ricci Curvature, and Einstein Metrics, geometry seminar at the Università di L’Aquila, L’Aquila, Italy.

31. May 2, 2003 – Transverse Fano Structures, and Einstein Metrics, geometry seminar at the University of California, Riverside, California, USA.

32. April 30, 2003 – Transverse Fano Structures and Positive Ricci Curvature, colloquium at the University of California, Riverside, California, USA.

33. April 8, 2003 – Transverse Fano Structures, Ricci Curvature, and Einstein Metrics, colloquium at the Oregon State University, Corvalis, Oregon, USA.

34. April 8, 2003 – Transverse Fano Structures, Ricci Curvature, and Einstein Metrics, geometry seminar at the University of Oregon, Eugene, Oregon, USA.

35. November 22, 2002 – Transverse Fano Structures, Positive Ricci Curvature, and Einstein Metrics, invited lecture at the Erwin Schrödinger Institute, Vienna, Austria.

36. June 15, 2002 – On the Geometry of Sasakian-Einstein 5-Manifolds, invited lecture at the Joint International AMS/UMI Meeting in Pisa, June 12-16, Pisa, Italy.

37. June 6, 2002 – Sasakian Manifolds and Einstein Metrics in Dimension 5, invited lecture at the EDGE meeting on Special and Low-dimensional Geometries, June 3-7, Luminy, France.

38. May 28, 2002 – Contact Structures, Einstein Metrics and Positive Ricci Curvature, invited lecture at Institut Élie Cartan, Université Henri Poincaré, Nancy, France.

39. May 27, 2002 – Sasakian Manifolds and Einstein Metrics, Arthur Besse seminar, École Polytechnique, Palaiseau, France.

40. December 6, 2001 – Metric Cones, Holonomy, and Special Metrics, the SFB 288 Differentialgeometrie und Quantenphysik Kolloquium, Freie Universität Berlin, Berlin, Germany.

41. November 24, 2001 – Contact Structures and Algebraic Geometry, invited lecture at the conference Recent trends in mathematics, November 23-25, 2001, Schloss Klink, Germany.

42. November 19, 2001 – On Sasakian-Einstein 5-Manifolds, Humboldt Universität geometry seminar, Berlin, Germany.

43. September 26, 2001 – New Einstein Metrics in Dimensions Five and Seven, lecture given at the Università di Roma, “La Sapienza”, Rome, Italy.

44. August 4, 2001 – 3-Sasakian Geometry, Nilpotent Orbits, and Exceptional Quotients, LMS Durham Symposium, Special Structures in Differential Geometry, University of Durham, July 30 – August 9, 2001, Durham, England.

45. June 11, 2001 – Algebraic Geometry and Einstein manifolds in Dimension Five, invited lecture at Mathematisches Forschunginstitut Oberwolfach, Differentialgeometrie im Grossen, June 11 – 15, Oberwolfach, Germany.

46. May 22, 2001 – Algebraic Geometry and Compact Einstein 5-Manifolds, lecture given at the geometry seminar at the University of Bonn, Bonn, Germany.

47. March 21, 2001 – Sasakian-Einstein Geometry of 5-Manifolds, lecture given at the Università di Roma, “La Sapienza”, Rome, Italy.

48. May 19, 2000 – New Einstein Metrics on 5-Manifolds, Arthur Besse seminar, Ecole Polytechnique, Palaiseau, France.

49. August 29, 1999 – Killing Spinors, Einstein Manifolds, and Contact Geometry, invited lecture at the conference on Generalized Dirac Operators and Their Geometric Scattering Theory, Warsaw, 23 August - 4 September, 1999; International Stefan Banach Center of the Polish Academy of Sciences, Warsaw, Poland.

50. July 13, 1999 – Sasakian-Einstein Manifolds, Killing Spinors, and Near Horizon Geometries, Humboldt Universität geometry seminar, Berlin, Germany.

51. June 17, 1999 – Contact Structures, Einstein Manifolds, and Near Horizon Geometries, Universität München geometry seminar, München, Germany.

52. February 24, 1999 – From Archimedes to Newton – A Beginner’s Guide to Symplectic Reduction, lecture at the West Washington University, Bellingham, Washington, USA.

53. February 22, 1999 – Cones, Holonomy, and Einstein Manifolds colloquium talk at West Washington University, Bellingham, Washington, USA.

54. June 18, 1998 – New Families of Einstein 7-manifolds, colloquium talk at Università di Bari, Bari, Italy.

55. June 2,4,9,2, 1998 – A Short Introduction to 3-Sasakian Manifolds I-IV, four lectures given at the Università di Roma, “La Sapienza”, Rome, Italy.

56. December 4, 1997 – Quaternionic Geometry and Einstein Manifolds, colloquium given in the University of Leipzig, Leipzig, Germany.

57. November 9, 1997 – Hypercomplex Structures on Toral Quotients, talk given in the Special Session of the AMS Meeting, Albuquerque, New Mexico, USA.

58. October 9, 1997 – Einstein Manifolds From Spheres, report on my research presented to the Fachbeirat at the Max-Planck-Institut für Mathematik in Bonn, Bonn, Germany.

59. August 18, 1997 – Quaternionic Geometry and Einstein Manifolds, lecture at the ICMS conference Complex Methods in Differential Geometry 18-29 August, 1997, Edinburgh, Scotland.

60. July 15, 1997 – Quaternionic Geometry and Einstein Manifolds, lecture given in the MPI Faltings-Harder-Manin-Zagier Oberseminar, MPI, Bonn, Germany.

61. June 9, 1997 – Making Einstein Manifolds Out of Spheres, geometry seminar at the Jagiellonian University, Kraków, Poland.

62. April 7, 1997 – On Betti Numbers of 3-Sasakian Manifolds, Topologie Oberseminar MPI, Bonn, Germany.

63. February 25,27 1997 – Compact Einstein 7-manifolds of Positive Scalar Curvature and Arbitrary Second Betti Number, and 3-Sasakian Geometry and Killing Spinors, two lectures at the Meeting on Dirac Operators and Geometry, Warsaw, 16 February – 1 March, 1997, International Stefan Banach Center of the Polish Academy of Sciences, Warsaw, Poland.

64. February 4, 1997 – Compact Einstein 7-manifolds of Positive Scalar Curvature and Arbitrary Second Betti Number, geometry seminar at the University of Bonn, Bonn, Germany.

65. November 13, 1996 – Making Einstein Manifolds Out of Spheres, colloquium at the University of California at Riverside, Riverside, California, USA.

66. June 4, 1996 – Geometry and Topology of 3-Sasakian Manifolds, I-III, series of three lectures at the ICMS Workshop Interplays Between Geometry and Topology, June 3-29, 1996, Edinburgh, Scotland.

67. April 14, 1996 – Hypercomplex manifolds and 3-Sasakian geometry, contributed paper in a Special Session at the AMS meeting, April 13-14, 1996, New York, New York, USA.

68. December 1, 1995 – On Betti Numbers of 3-Sasakian Manifolds, contributed paper in a Special Session at the second joint meeting of AMS and SMM, November 29 – December 2, 1995, Guanajuato, Mexico.

69. November 21, 1995 – Geometry and Topology of 3-Sasakian Manifolds: On Betti Numbers of 3-Sasakian Spaces, Humboldt Universität geometry seminar, Berlin, Germany.

70. November 16, 1995 – Quaternionic Geometry, Supergravity, and Einstein Manifolds, colloquium at Augsburg University, Augsburg, Germany.

71. November 14, 1995 – Geometry and Topology of 3-Sasakian Manifolds: Quotient Construction, Humboldt Universität geometry seminar, Berlin, Germany.

72. November 9, 1995 – Quaternionic Geometry, 3-Sasakian Manifolds, and Supergravity, the SFB 288 Differentialgeometrie und Quantenphysik Kolloquium, Freie Universität Berlin, Berlin, Germany.

73. June 12, 1995 – 3-Sasakian Manifolds, invited lecture at Mathematisches Forschunginstitut Oberwolfach, Differentialgeometrie im Grossen, June 12 – 16, Oberwolfach, Germany.

74. June 9, 1995 – Geometry and Topology of 3-Sasakian Manifolds, geometry seminar at the Max-Planck-Institut, Bonn, Germany.

75. May 22, 1995 – Geometry and Topology of 3-Sasakian Manifolds, invited lecture in Arthur Besse Seminar at École Polytechnique, Paris, France.

76. January 10, 1995 – 3-Sasakian Manifolds, lecture presented at the conference on Representation Theory and Differential Geometry, University of Adelaide, Adelaide, Australia.

77. September 28, 1994 – Einstein Metrics and 3-Sasakian Geometry, lecture presented at Erwin Schrödinger Institute, Vienna, Austria.

78. September 9, 1994 – On 3-Sasakian manifolds, lecture presented at the Meeting on Quaternionic Structures in Mathematical Physics, September 5 – 9, 1994; SISSA, Trieste, Italy.

79. November 17, 1993 – On strongly inhomogeneous Einstein manifolds with positive scalar curvature and 3-Sasakian structure, talk presented at Columbia University, New York, New York, USA.

80. November 16, 1993 – On strongly inhomogeneous Einstein manifolds with positive scalar curvature and 3-Sasakian structure, talk presented at SUNY at Stony, Brook, Stony Brook, New York, USA.

81. November 15, 1993 – On strongly inhomogeneous Einstein manifolds with positive scalar curvature and 3-Sasakian structure, talk presented at the Institute for Advanced Study, Princeton, New Jersey, USA.

82. July 1, 1993 – A construction of strongly inhomogeneous Einstein manifolds with 3-Sasakian structure, talk presented at the Max-Planck Institut Harder-Hirzebruch-Zagier Oberseminar, Max-Planck Institut für Mathematik, Bonn, Germany.

83. June 4, 1993 – Einstein manifolds in field theories with supersymmetry (in Polish), talk given at the University of Wrocław, Wrocław, Poland.

84. May 26, 1992 – O pewnych siedmiowymiarowych rozmaitościach Einsteina (in Polish), Wrocław University, Wrocław, Poland.

85. February 29, 1992 – Quaternionic Reduction and Einstein Manifolds, The New Mexico Geometry and Topology Symposium, Las Cruces, New Mexico, USA.

86. February 12, 1991 – Group Actions, Quotients, and Quaternionic Manifolds, invited talk at SUNY@Stony Brook, Stony Brook, New York, USA.

87. January 27, 1991 – Group Actions, Quotients, and Quaternionic Manifolds, talk given at The Institute for Advanced Study, Princeton, New Jersey, USA.

88. May 14, 1990 – Self-dual Einstein Metrics with Nonzero Scalar Curvature, talk given at The University of Warwick, England.

89. February 6, 1990 – Quaternionic Kähler Reduction and Self-Dual Einstein Orbifolds, Washington University, St. Louis, Missouri, USA.

90. February 1, 1990 – Quaternionic Kähler Reduction and Self-Dual Einstein Orbifolds, The University of New Mexico, Albuquerque, New Mexico, USA.

91. January 29, 1990 – Quaternionic Kähler Reduction and Self-Dual Einstein Orbifolds, University of British Columbia, Vancouver, BC, Canada.

92. January 28, 1990 – Quaternionic Kähler Reduction and Self-Dual Einstein Orbifolds, University of California at Davis, Davis, California, USA.

93. January 25, 1990 – Quaternionic Kähler Reduction and Self-Dual Einstein Orbifolds, Oregon State University, Corvalis, Oregon, USA.

94. January 23, 1990 – Quaternionic Kähler Reduction and Self-Dual Einstein Orbifolds, Oklahoma State University, Stillwater, Oklahoma, USA.

95. January 11, 1990 – Quaternionic Kähler Reduction and Self-Dual Einstein Orbifolds, SUNY@Buffalo, Buffalo, New York, USA.

96. August 25, 1989 – Quotient Constructions of Special Riemannian Geometries, Odense University, Odense, Denmark.

97. August 23, 1989 – Quotient Constructions of Special Riemannian Geometries, Danmarks Tekniske Høskole, Copenhagen, Denmark.

98. June 16, 1989 – Quotient Constructions of Special Riemannian Geometries, Max-Planck-Institut für Mathematik, Bonn, Germany.

99. February 21, 1989 – Generalized Symplectic Reduction, University of California, Santa Barbara, California, USA.

100. December 8, 1988 – New metrics with Sp(n) · Sp(1) holonomy, Tulane University, New Orleans, Louisiana, USA.

101. July 6-15, 1988 – Quaternionic reduction and quaternionic orbifolds, The Second Meeting on Applications of Group Theory to Math-Physics, University of Georgia, Athens, Georgia, USA.

102. February 27, 1988 – Quaternionic reduction and quaternionic orbifolds, SUNY at Stony Brook, Stony Brook, New York, USA.

103. October 12, 1987 – A generalization of moment map in quaternionic geometry, Texas A&M University, College Station, Texas, USA.

104. April 28, 1985 – Quaternionic manifolds and local supersymmetry, University of Georgia, Athens, Georgia, USA.

Last updated: February 1, 2006